It's all about excel!

  1. How many new clients did Vantage Investments gain in 2016-17? What is the total sales volume ($)? On average, how much new investment did each new client bring in?
  2. What is the average number of new clients brought in by a sales person? What is the average value of the new investments brought in by a salesperson? What is the standard deviation of the new investments brought in by a salesperson?
SALESPERSON 2016 SALES TARGET 2016 NEW CLIENTS 2016 NEW SALES
Almeida 215000 15 212723
Anderson 170000 16 171616
Ball 155000 13 153201
Ballestero 125000 12 124904
Barrio 215000 15 213756
Billimoria 65000 5 66519
Budge 185000 14 183450
Cohen 95000 10 96370
Curtin 185000 15 183008
Drew 185000 12 186783
Answer #1

Solution:

Calculation of Total new clients & Total new sales in 2016

SALESPERSON 2016 SALES TARGET 2016 NEW CLIENTS 2016 NEW SALES
Almeida 215000 15 212723
Almeida 170000 16 171616
Ball 155000 13 153201
Ballestero 125000 12 124904
Barrio 215000 15 213756
Billimoria 65000 5 66519
Budge 185000 14 183450
Cohen 95000 10 96370
Curtin 185000 15 183008
Drew 185000 12 186783
TOTAL 127 1592330
1) Vantage investments gain 127 new clients in 2016.
2) The total sales volume is $1592330.
3) Avg investments brought by each client = Total new sales / Number new clients
Avg investments brought by each client = $1592330 / 127 = $12538
4) Average number of new client brought in by a sales person
= Total new clients / Total Number of sales persons = 127 / 10 = 12.7
5) Average value of the new investments brought in by a salesperson
= Total new sales / Total Number of sales person
= $1592330 / 10 = $159233